OpenStudy (anonymous):

Based on the numbers in the previous question, an 80–pound Earth girl would weigh about ___ pounds on the planet Namar. A. 4 B. 320 C. 18 D. 40

OpenStudy (lifeisadangerousgame):

What was the previous question?

OpenStudy (anonymous):

18/4=80/X X=80*4/18